Grade 12th passMechanics

A 5 × 105 kg rocket is accelerating straight up. Its engines produce 1.25 × 107
N of thrust, and air resistance is 4.5 × 106 N. What is the rocket’s acceleration?

hilal shah
6 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer6 Years ago

Since the rocket is moving upward, its weight is causing it to decelerate.

Weight = 5 * 10^5 * 9.8 = 4.9 * 10^6 N

The net force is equal to the thrust force minus the sum of the weight and the air resistance.

 

Net force = 1.25 * 10^7 – (4.5 * 10^6 + 4.9 * 10^6) = 3.1 * 10^6 N

F = m * a

3.1* 10^6 = 5 * 10^5 * a

a = 3.1* 10^6 ÷ 5 * 10^5 = 6.2 m/s^2